What is the molarity of solution prepared by dissolving 11.75 g of KNO3 in enough water to produce 2.00 L of solution?
Answer #1

M = mol /L

m = 11.75 g of KNO3

MW = 101.1032

mol = mass/MW = 11.75/101.1032 = 0.1162178

V = 2 L

Then

M = 0.1162178/2 = 0.0581 M